OpenStudy (anonymous):

What is the solution to the system? Line 1 x y –1 –2 1 2 Line 2 x y –2 10 4 1 ( , ) please help will fan and medal.

OpenStudy (bohotness):

Line 1: (-1,-2) and (1,2) Line 2: (-2,10) and (4,1) ---- Get equations of the lines. Line 1: (-1,-2) and (1,2) Find the slope, m m = diffy/diffx = (-2-2)/(-1-1) = 2 Use y = mx + b and either point to find b, the y-intercept Using (1,2): 2 = 2*1 + b --> b = 0 y = 2x is an equation of Line 1 ---- Get the eqn of line 2 the same way.

OpenStudy (anonymous):

im confused

OpenStudy (bohotness):

okay

OpenStudy (bohotness):

let me try this one

OpenStudy (bohotness):

what is the context of this ?

OpenStudy (bohotness):

is it a linear equation?

OpenStudy (bohotness):

@bcimkenz is it a linear equation?